Home‎ > ‎Crew 162‎ > ‎

0129-DailySummaryReport

Crew 162 Daily Summary Report for Sol 05 01/29/2016

Summary Title: Rest, Relaxation and Research

Mission Status: Sol 06 served as a rest and recovery day, providing the
crew an opportunity to manage their own individual schedules and perform
checks on equipment and experiments with ample free time to rest after
yesterday's exhausting timetable. Crew is ready to take on the second
half of our mission.

Sol Activity Summary: Crew 162 gathered for a late breakfast and
planning meeting. Individual research and rest were the theme for today.
Commander worked on follow-up activities, greenhab experiments, soil
analysis, video and photo editing. Health & Safety Officer worked on
blogs, editing videos for her Spanish-language documentary and
troubleshooting the 3-D printer. Crew Engineer performed engineering
checks and documented ATV wheel status for engineering support.
Commander assisted with the engineering checks and helped document video
of the ATV checks. We had a much-needed relaxed lunch a Mexican chicken
burrito theme. We discussed our experiences at ISU, our graduate school,
and shared stories about family. We are nearing the half-way point of
our mission and the hab is feeling more like home here at the Martian
outpost.

Look Ahead Plan: Bad weather predicted this weekend, possible snow on
Sunday Sol 08. We will have our second media visit on Tuesday, Sol 10,
by a professional photographer from London.

Anomalies in work: No current anomalies to report. We are currently
locked out of the observatory but astronomy mission support is working
to get a replacement padlock and key.

Weather: Sunny and pleasant day on Sol 06. Low humidity and very little
wind.

Crew Physical Status: Rested and well-fed. Crew is feeling positive
about the second half of our mission.

EVA: No EVA today, only engineering checks by Crew Engineer accompanied
by Commander.

Reports to be filed:
Commanders Report
Engineering Report
Journalist Report
GreenHab Report
EVA #3 Request
Daily Photos

Support Requested: No support requested from Mission Control at this
time.